OpenStudy (anonymous):

compute the descriminant of the quadractic equation 2x^2 -3x-5 = 0 then descriding the number and type of solutions for the equation. please help this is my answer x1 5/2 but i need what they are asking for

myininaya (myininaya):

do b^2-4ac

myininaya (myininaya):

=3^2-4(2)(-5) so this is positive two real solutions

OpenStudy (anonymous):

the general form is ax^2 + bx +c= 0 2x^2 -3x-5 = 0 so a=2 b=-3 and c =-5

OpenStudy (anonymous):

thank you myinnaya and dindatc